When nitrogen dioxide (NO2) from car exhaust combines with water in the air, it forms nitric acid (HNO3), which causes acid rain, and nitrogen oxide.

3NO2(g)+H2O(l)→2HNO3(aq)+NO(g)

A. How many moles of HNO3 are produced from 0.196 mole of H2O?

B.How many moles of NO are produced from 0.196 mole of H2O?

C.How many grams of HNO3 are produced when 90.5 g of NO2 completely reacts?

D.How many grams of NO2 are needed to form 64.5 g of HNO3?

Gasohol is a fuel containing ethanol (C2H6O) that burns in oxygen (O2) to give carbon dioxide and water

.A. What is the chemical equation with phases

B. If a car produces 73 g of CO2, how many grams of O2 are used up in the reaction?

C. If you burn 136 g of C2H6O, how many grams of CO2 and H2O can be produced?
Answer #1

3NO2(g)+H2O(l)→2HNO3(aq)+NO(g)

0.196 mole of H2O will produce = 0.196 x 2 = 0.392 Mole HNO3 will produce.

0.196 mole of H2O will produce = 0.196 x 1 = 0.196 Mole NO will produce

90.5 g of NO2 = 90.5 /46 = 1.96 Mole x 63.01 x 2 /3 = 82.64 gm of HNO3 will produce.

64.5 g of HNO3 = 64.5 / 63.01 = 1.0236 x 46 x 3/2 = 70.63 gm of NO2 is need

C + O2 = CO2

73 g of CO2 = 73/44 = 1.659 Mole x 32 = 53.09 gm of O2 were used up.

C2H6O + 3 O2 = 2 CO2 + 3 H2O

136 g of C2H6O = 136 / 46.07 = 2.95 Mole x 2 x 44 = 259.77 gm of CO2 and

                                136 / 46.07 = 2.95 Mole x 3 x 18 = 159.3 gm of H2O will produce